# The inclusive Synthetic Control Method

### v1

@@ Abstract
We introduce the inclusive synthetic control method (iSCM), a modification of synthetic control type methods that allows the inclusion of units potentially affected directly or indirectly by an intervention in the donor pool. This method is well suited for applications with either multiple treated units or in which some of the units in the donor pool might be affected by spillover effects. Our iSCM is very easy to implement using most synthetic control type estimators. As an illustrative empirical example, we re-estimate the causal effect of German reunification on GDP per capita allowing for spillover effects from West Germany to Austria.
Keywords: Synthetic Control Method, spillover effects, causal inference.
@@ 1 Introduction
The synthetic control method (SCM) introduced by Abadie and Gardeazabal (2003) and further developed in Abadie et al. (2010) and Abadie et al. (2015) allows the estimation of the causal effect of a policy intervention in settings where only a few treated and control units are observed over a long time period. The idea behind this method is to create a linear combination of control units that mimic what would have happened to the treated units in the absence of the intervention. The weights given to each control unit are chosen to minimize the distance in the pre-intervention characteristics of the treated and the synthetic control units. The causal effect of the intervention is then estimated as the difference between the observed outcome of the treated and that of the synthetic control in the post-intervention period. One of the key assumptions of the SCM is that only units that are not affected by the interventions are included in the control group, often referred to as the donor pool. This might be problematic in at least two scenarios: i) some of the treated units would ideally be included in the donor pool to improve the pre-intervention fit; ii) some of the control units are affected by the intervention indirectly but excluding them from the donor pool deteriorates the pre-intervention fit substantially. As a motivating example, consider the German reunification study by Abadie et al. (2015) (see also Abadie (2021)). The authors point out it is possible that German reunification had spillover effects on neighboring countries such as Austria. As Austria receives a high weight (42%), such a spillover effect, if large, would introduce a large bias. Given that Austria plays an important role in constructing “synthetic West Germany” excluding it from the donor pool is likely to reduce the quality of the match between treated and synthetic units.
Our main contribution is to introduce the inclusive synthetic control method (iSCM), a novel procedure that allows us to eliminate post-intervention effects from control units and safely include them in the donor pool. We exploit the fact that the synthetic control weights are estimated only using pre-intervention data and no unit is treated in this period. Thus, when “potentially affected” units are added to the donor pool, the synthetic control unit becomes a weighted average of the outcomes of all the units in the donor pool and the effects of the intervention on “potentially affected” units, each weighted by their synthetic weights. This implies that comparing the “main treated” and its synthetic control version gives an estimate of the true effect minus the effects on other units, each weighted by their synthetic weights. If we create a synthetic control version of all “potentially affected” units (including the “main treated” unit in the respective donor pools) we can get an estimate of the true effects on those units plus the effects on other affected units (including the “main treated”), each weighted by their synthetic weights. This process yields a system of equations with “m” unknowns, comprising the treatment effect on the “main treated” unit and the “m-1” effects on “potentially affected” units which can be easily solved using Cramer’s rule. This means that our procedure does not require any modification of the synthetic control estimator, and any alternative SC-type method which creates a synthetic control unit as a weighted average of the outcome of the units in the donor pool can be used instead (see, for example, Abadie and L’Hour 2021, Amjad et al. 2018, Ben-Michael et al. 2021, Ben-Michael et al. 2022, Doudchenko and Imbens 2017, Ferman and Pinto 2021, Kellogg et al. 2021, Xu 2017). In addition to the assumptions behind the particular SC-estimator chosen, our iSCM requires prior knowledge on which units can be “potentially affected” by the treatment. Although iSCM only requires the existence of at least one “pure control” unit receiving non-zero weight, our method becomes less reliable as the number of “potentially affected” units increases. Thus, it is advisable to impose assumptions that limit this number. This is in line with the literature on spillover effects, in which it is often assumed that interactions between units are only possible if they belong to a certain group, but no interactions are allowed between different groups (Cerqua and Pellegrini 2017, Forastiere et al. 2021, Huber and Steinmayr 2019, Vazquez-Bare 2022). This is also the case for the few extensions of the synthetic control method that allow for the presence of spillover effects. Grossi et al. (2020) propose reducing the donor pool to only units not affected by spillovers. They estimate the effects for the treated unit using a SCM with the restricted donor pool and the spillover effects by comparing units affected by spillover and the restricted donor pool. Their method is very effective in applications in which the restricted donor pool is sufficient to construct a “good” synthetic control. However, in a setting where including units affected by spillover in the donor pool improves the pre-intervention fit, as in the German reunification example, their method would likely produce less reliable results. Cao and Dowd (2019) provide a different identification and estimation strategy focusing more on the spillover effects structure. In contrast, our approach can also be used in applications where one wants to include treated units in the donor pool. Another difference between the two approaches is that we allow for using any SC-type estimator as soon as the final estimator is based on a weighted average of the outcome of the units in the donor pool. Although our paper is related to the literature on spillover effects, our results also apply when there are multiple treated units and no spillover effects. Moreover, our identification relies on being able to observe the potential outcome under non-treatment of every unit in the pre-intervention period and it differs from the identification results in the spillover/peer effect literature.
We provide a data-driven approach to check whether excluding “potentially affected” units from the donor pool can potentially lead to a large bias. Even in applications where the expected bias is small we still recommend using iSCM as a robustness check especially if the “potentially affected” units receive a high weight.
@@
The rest of the paper is organized as follows: Sections 2 describe the iSCM; Section 3 discuss the special case where there is only one “potentially affected” unit; Section 4 compare the estimation errors of SCM and iSCM; in Section 5 we compare iSCM and a “restricted” SCM and propose an implementation procedure; Section 6 suggests possible inference procedures; Section 7 shows the results of the empirical application; and Section 8 concludes the paper.
@@ 2 The inclusive synthetic control method
Assume we observe $j=1,...,J$ units for $t=1,...,T$ periods. For each unit $j$ at time $t$ we observe the outcome of interest $Y_{jt}$ and a set of $k$ predictors of the outcome $X_{1j},...,X_{kj}$, which often includes pre-intervention values of $Y_{jt}$. We refer to the treated unit (unit $1$) as the “main treated”. We also assume that the $J-1$ units in the donor pool, include $m<J-1$ units (units $2$ to $m$) that are directly or indirectly affected by the intervention (“potentially affected” hereafter), i.e., they are either other treated units that we would like to include in the donor pool or control units that might be affected by spillover effects from the main treated. We refer to units $m+1$ through $J$, as “pure control” units and assume that they are not affected by the intervention at all. We define the potential outcome $Y_{1t}^{I}$ as the outcome that the “main treated” unit would obtain under the intervention at time $t$. We define the potential outcome of “potentially affected” units $j$ at time $t$ as $Y_{jt}^{S},j=2,\ldots,m$, which represents either the outcome of a unit “potentially affected” by spillover effects or simply the potential outcome under treatment if unit $j$ is a different treated unit. Finally, we define $Y_{jt}^{N},j=1,\ldots,J$ as the potential outcome in the absence of the intervention. We assume that there are no anticipation effects together with the standard stable unit treatment value assumption (SUTVA) with the exception of the potential presence of spillover effects on units, who are chosen a priori. This assumption can be formalized as Assumption 1:
@@
  • In the post-intervention period, $Y_{jt}$=$Y_{jt}^{N}$ for the “pure control”.
  • In the post-intervention period, $Y_{1t}$=$Y_{1t}^{I}$ for the “main treated”.
@@
Assumption 1 may be restrictive in scenarios where it is unclear which units might be “potentially affected” by spillover effects. However, it is standard in contexts where the “potentially affected” units are other treated units. We are interested in the causal effect of the intervention on the “main treated” at time $t>T_{0}$, denoted by $\theta_{1t}$, and the causal effects on the other “potentially affected” units denoted by $\gamma_{jt},j=2,\ldots,m,t>T_{0}$, defined as
@@
Assume we use a SC-type estimator $\widehat{Y}_{1t}^{N}=\sum_{j=2}^{J}\widehat{w}_{j}Y_{jt}$, which estimates $Y_{1t}^{N}$ as a weighted average of the post-intervention outcomes of the units in the donor pool, e.g., the original SCM as described in Abadie et al. (2010). Let the $\left((J-1)\times 1\right)$ vector of weights $\widehat{W}=(\widehat{w}_{2},\ldots,\widehat{w}_{J})^{\prime}$ be the estimated weights of the chosen SC-type estimator that includes also the “potentially affected” units in the donor pool, under Assumption 1, $\forall\ t>T_{0},$ we have
@@
Similarly, consider a generic “potentially affected” unit $i$, $i\in M\equiv[2,\ldots,m]$. Let $\widehat{L}^{i}$ the vector of weights estimated to create $\widehat{Y}_{it}^{N}=\sum_{j\neq i}\widehat{l}_{j}^{i}Y_{jt}$ a SC-type estimator of $Y_{it}^{N}$ that includes the “main treated” (unit 1), the other $m-1$ “potentially affected” units, as well as the $J-m$ “pure” control units in the donor pool. Assumption 1 implies that, $\forall\ i=2,\ldots,m,\ t>T_{0}$, we have
@@
$=$ $\sum_{j=m+1}^{J}\widehat{l}_{j}^{i}Y^{N}_{jt}+\sum_{j\in M\backslash\{{i}\}}\widehat{l}_{j}^{i}\left(\underbrace{Y^{N}_{jt}+\gamma_{jt}}_{Y^{S}_{jt}}\right)+\widehat{l}^{i}_{1}{\left(\underbrace{Y^{N}_{1t}+\theta_{jt}}_{Y^{I}_{jt}}\right)},$
@@
Looking at the bias decomposition in (1) and (2) it is clear that we have two distinct sources of bias. The first, $B^{j}_{sc},j=1,\ldots,m$, relates to how well the chosen SC-type estimators approximate the post-intervention counterfactual. This is the usual source of bias we encounter when using any SC-type estimator. The second, $B^{j}_{te},j=1,\ldots,m$, is induced by the fact that we are using units that are “potentially affected” by the intervention in the donor pull. This type of bias solely depends on whether or not a given (“potentially affected”) unit receives weight and its respective treatment effect but not on how “good” the chosen estimator is in recovering $Y_{jt}^{N},j=1,\ldots,m$.
@@
Assumption 2: The SC-type used to estimate $\widehat{W}$ satisfy $\sum_{j=2}^{J}\widehat{w}_{j}{Y_{jt}^{N}}=Y_{1t}^{N}+o_{p}(1)$.
We state Assumption 2 in terms of estimated weights without any loss of generality and simplify the discussion. Alternatively we can assume that the set of weights $\widehat{W}$ converges to a set of weights $W^{*}$ such that $\sum_{j=2}^{J}w^{*}_{j}{Y_{jt}^{N}}=Y_{1t}^{N}+o_{p}(1)$. When using the original SC-estimator, Assumption 2 amounts to assuming an (approximately) perfect fit444see the discussion in Abadie (2021), Ferman and Pinto (2021), and Powell (2018) about the lack of a perfect fit.. We state assumption 2 in very general terms as our results apply to any SC-type estimator that can be written as a weighted average of the control units post-intervention outcomes. For example, for the original SCM one can use the results in Zhang et al. (2022), and state Assumption 2 accordingly. Depending on the chosen estimator, one can state Assumption 2 more precisely. A non-exhaustive list of possible estimators includes the one proposed in: Abadie and L’Hour 2021, Amjad et al. 2018, Ben-Michael et al. 2021, Ben-Michael et al. 2022, Doudchenko and Imbens 2017, Ferman and Pinto 2021, Kellogg et al. 2021, Xu 2017.
@@
It is important to notice that for each unit $j=2,\ldots,m$, if either $\gamma_{jt}$ or $\widehat{w}_{j}$ is zero, that unit does not induce any extra bias in $\widehat{\theta}_{1t}$. This implies that units that receive a low estimated weight need to have an extremely large effect to induce a non-negligible bias in $\widehat{\theta}_{1t}$. For this reason, units that receive a low weight can be relatively safely treated as “pure controls” when estimating $\theta_{1t}$ in empirical applications.
@@
When using SC-type estimators that allow for negative weights one needs to check the magnitude of the weights given to “potentially affected” units. For estimators that use a bias correction, one needs to subtract the estimated bias from the outcome of all units before running our iSCM (see equation (16) in Abadie 2021).
@@
Assumption 3: The SC-estimator chosen to estimate $\widehat{L}^{i}$ satisfy $\widehat{Y}_{it}^{N}=Y_{it}^{N}+o_{p}(1),\ \forall\ i=2,\ldots,m.$
@@
This is a system of $m$ equations with $m$ unknowns, i.e., the treatment effect on the “main treated” and the $m-1$ effects on the “potentially affected” units.
@@
It is easy to show that $\widehat{\Omega}$ is always invertible, if $m<J-1$, except for the extreme cases in which two units give weight 1 to each other and/or every single weight associated with the “pure control” units is zero (see Appendix A.1).
@@
3 An example with a single “potentially affected” unit
To further illustrate our results, it is useful to consider the special case in which, together with the “main treated” unit, only one additional unit is “potentially affected” by the intervention ($m=1$). First, we show the standard case and then the simplified case in which including the “main treated” in the donor pool of the “potentially affected” unit is not necessary.
With only one “potentially affected” unit, the system of equation defined in Section 2, again ignoring the estimation errors, simplifies to

### v2

@@ Abstract
We introduce the inclusive synthetic control method (iSCM), a modification of synthetic control methods that includes units in the donor pool potentially affected, directly or indirectly, by an intervention. This method is ideal for situations where including treated units in the donor pool is essential or where donor units may experience spillover effects. The iSCM is straightforward to implement with most synthetic control estimators. As an empirical illustration, we re-estimate the causal effect of German reunification on GDP per capita, accounting for spillover effects from West Germany to Austria.
Keywords: Causal inference, iSCM, Multiple treated, Spillover effects, Synthetic Control Method.
@@ 1 Introduction
The synthetic control method (SCM), introduced by Abadie and Gardeazabal (2003) and further developed in Abadie et al. (2010) and Abadie et al. (2015), estimates the causal effect of a policy intervention in settings with few treated and control units observed over a long time period. SCM constructs a linear combination of control units to mimic what would have happened to the treated units had the intervention not occurred. The weights for each control unit are chosen to minimize differences in pre-intervention characteristics between the treated units and the synthetic control. The causal effect is then estimated as the difference between the observed outcome of the treated units and that of the synthetic control in the post-intervention period.
A key assumption of SCM is that only units not affected by the intervention are included in the control group, often called the donor pool. This can be problematic in scenarios where (i) some treated units should ideally be included in the donor pool to improve pre-intervention fit, or (ii) some control units are indirectly affected by the intervention, but excluding them from the donor pool substantially deteriorates the pre-intervention fit.
As a motivating example, consider the German reunification study by Abadie et al. (2015) (see also Abadie (2021)). The authors suggest that reunification may have had spillover effects on neighboring countries like Austria. Since Austria receives a substantial weight (42%) in constructing ”synthetic West Germany,” any significant spillover effect could introduce considerable bias. Excluding Austria from the donor pool substantially reduces the quality of the match between the treated and synthetic units. This is also the case when using the penalized SCM (see, Section 7.1). Moreover, excluding West Germany from Austria’s donor pool when estimating the spillover effect yields implausible positive spillover effects on Austria’s GDP.
Our main contribution is the introduction of the inclusive synthetic control method (iSCM), a novel procedure that allows us to eliminate post-intervention effects from control units and safely include them in the donor pool. We exploit the fact that synthetic control weights are estimated using only pre-intervention data, during which no unit is treated. By adding ”potentially affected” units to the donor pool, the synthetic control becomes a weighted average of all donor units’ outcomes, including any intervention effects on these units, weighted by their synthetic weights. Comparing the ”main treated” unit and its synthetic control provides an estimate of the true effect minus the effects on other units, each weighted by their synthetic weights.
By creating synthetic control versions of all ”potentially affected” units (including the ”main treated” unit in their respective donor pools), we obtain estimates of the true effects on those units plus the effects on other affected units, weighted by their synthetic weights. This process yields a system of equations with ”m” unknowns—the treatment effect on the ”main treated” unit and the ”m-1” effects on the ”potentially affected” units—which can be solved using Cramer’s rule. Importantly, our procedure does not require any modification of the synthetic control estimator. Any SCM-type method that creates a synthetic control as a weighted average of donor units’ outcomes can be used (e.g., Abadie and L’Hour 2021, Amjad et al. 2018, Ben-Michael et al. 2021, Ben-Michael et al. 2022, Doudchenko and Imbens 2017, Ferman and Pinto 2021, Kellogg et al. 2021, Xu 2017). In addition to the assumptions of the chosen SCM estimator, our iSCM requires prior knowledge of which units may be ”potentially affected” by the treatment.
While iSCM requires at least one ”pure control” unit receiving non-zero weight, the method becomes less reliable as the number of ”potentially affected” units increases. Therefore, it is advisable to impose assumptions that limit this number, aligning with the spillover effects literature, which often assumes interactions only within certain groups (Cerqua and Pellegrini 2017, Forastiere et al. 2021, Huber and Steinmayr 2019, Vazquez-Bare 2022). Existing extensions of SCM that account for spillover effects often reduce the donor pool to unaffected units. For instance, Grossi et al. (2024) estimate effects for the treated unit using SCM with a restricted donor pool and estimate spillover effects by comparing spillover-affected units with the restricted donor pool. Their method works well when the restricted donor pool generates a good synthetic control. However, when including spillover-affected units in the donor pool improves pre-intervention fit, as in the German reunification example, their method may be less reliable.
Cao and Dowd (2019) provide a different identification and estimation strategy focusing more on the spillover effects structure. In contrast, our approach can also be used in applications where one wants to include treated units in the donor pool. Another difference between the two approaches is that we allow for using any SC-type estimator as soon as the final estimator is based on a weighted average of the outcome of the units in the donor pool.
Although our paper is related to the literature on spillover effects, our results also apply when there are multiple treated units and no spillover effects. Moreover, our identification relies on being able to observe the potential outcome under non-treatment of every unit in the pre-intervention period and it differs from the identification results in the spillover/peer effect literature.
Our setting can also be viewed as a multiple treated units scenario, inviting a comparison to methods developed for such contexts (Abadie and L’Hour 2021, Ben-Michael et al. 2022, Kellogg et al. 2021). These methods either modify traditional SCM by introducing a penalty term to reduce large discrepancies between the treated unit’s pre-treatment characteristics and those of its synthetic control, or they combine SCM with matching estimators to balance interpolation and extrapolation biases. While these approaches are useful in high-dimensional or disaggregated data settings, they are not designed to handle scenarios where the ideal donor pool includes units potentially affected by the intervention, either directly or through spillover effects.
In contrast, our iSCM approach allows for the direct inclusion of these affected units in the donor pool without modifying the underlying SCM estimator (which could be any of the methods mentioned above). This is particularly beneficial in multiple treated units settings, where excluding treated or spillover-affected units could weaken the quality of the synthetic match, as demonstrated in our application.
In conclusion, our iSCM allows for the inclusion of units ”potentially affected” directly or indirectly by the intervention in the donor pool. Given its ease of implementation, iSCM can always be used, even as a robustness check. Additionally, we provide a data-driven procedure to determine whether iSCM is preferable to a restricted SCM that excludes these ”potentially affected” units from the donor pool.
@@
The rest of the paper is organized as follows: Sections 2 describe the iSCM; Section 3 discuss the special case where there is only one ”potentially affected” unit; Section 4 compare the estimation errors of SCM and iSCM; in Section 5 we compare iSCM and a ”restricted” SCM; Section 6 suggests possible inference procedures; Section 7 shows the results of the empirical application; and Section 8 concludes the paper.
@@ 2 The inclusive synthetic control method
Assume we observe $j=1,...,J$ units for $t=1,...,T$ periods and that an intervention occurs at time $T_{0}$. For each unit $j$ at time $t$ we observe the outcome of interest $Y_{jt}$ and a set of $k$ predictors of the outcome $X_{1j},...,X_{kj}$, which often includes pre-intervention values of $Y_{jt}$. We refer to the treated unit (unit $1$) as the ”main treated”. We also assume that the $J-1$ units in the donor pool, include $m<J-2$ units (units $2$ to $m$) that are directly or indirectly affected by the intervention (“potentially affected” hereafter), i.e., they are either other treated units that we would like to include in the donor pool or control units that might be affected by spillover effects from the main treated. We refer to units $m+1$ through $J$, as ”pure control” units and assume that they are not affected by the intervention at all. We define the potential outcome $Y_{1t}^{I}$ as the outcome that the ”main treated” unit would obtain under the intervention at time $t$. We define the potential outcome of ”potentially affected” units $j$ at time $t$ as $Y_{jt}^{S},j=2,\ldots,m$, which represents either the outcome of a unit ”potentially affected” by spillover effects or simply the potential outcome under treatment if unit $j$ is a different treated unit. Finally, we define $Y_{jt}^{N},j=1,\ldots,J$ as the potential outcome in the absence of the intervention. We assume there are no anticipation effects and that the standard Stable Unit Treatment Value Assumption (SUTVA) holds, except for the potential presence of spillover effects on units, who are chosen a priori. This assumption can be formalized as
Assumption 1:
@@
  • In the post-intervention period, $Y_{jt}$=$Y_{jt}^{N}$ for the ”pure control”.
  • In the post-intervention period, $Y_{1t}$=$Y_{1t}^{I}$ for the ”main treated”.
@@
Assumption 1 may be restrictive in scenarios where it is unclear which units might be ”potentially affected” by spillover effects. However, it is standard in contexts where the ”potentially affected” units are other treated units.
We are interested in the causal effect of the intervention on the ”main treated” at time $t>T_{0}$, denoted by $\theta_{1t}$, and the causal effects on the other ”potentially affected” units denoted by $\gamma_{jt},j=2,\ldots,m,t>T_{0}$, defined as
@@
Assume we use a SC-type estimator $\widehat{Y}_{1t}^{N}=\sum_{j=2}^{J}\widehat{w}_{j}Y_{jt}$, which estimates $Y_{1t}^{N}$ as a weighted average of the post-intervention outcomes of the units in the donor pool, e.g., the original SCM as described in Abadie et al. (2010). Let the $\left((J-1)\times 1\right)$ vector of weights $\widehat{W}=(\widehat{w}_{2},\ldots,\widehat{w}_{J})^{\prime}$ be the estimated weights of the chosen SC-type estimator that includes also the ”potentially affected” units in the donor pool, under Assumption 1, $\forall\ t>T_{0},$ we have
@@
Similarly, consider a generic ”potentially affected” unit $i$, $i\in M\equiv[2,\ldots,m]$. Let $\widehat{L}^{i}$ the vector of weights estimated to create $\widehat{Y}_{it}^{N}=\sum_{j\neq i}\widehat{l}_{j}^{i}Y_{jt}$ a SC-type estimator of $Y_{it}^{N}$ that includes the ”main treated” (unit 1), the other $m-1$ ”potentially affected” units, as well as the $J-m$ ”pure” control units in the donor pool. Assumption 1 implies that, $\forall\ i=2,\ldots,m,\ t>T_{0}$, we have
@@
$=$ $\sum_{j=m+1}^{J}\widehat{l}_{j}^{i}Y^{N}_{jt}+\sum_{j\in M\backslash\{{i}\}}\widehat{l}_{j}^{i}\left(\underbrace{Y^{N}_{jt}+\gamma_{jt}}_{Y^{S}_{jt}}\right)+\widehat{l}^{i}_{1}{\left(\underbrace{Y^{N}_{1t}+\theta_{1t}}_{Y^{I}_{jt}}\right)},$
@@
Looking at the bias decomposition in (1) and (2) it is clear that we have two distinct sources of bias. The first, $B^{j}_{sc},j=1,\ldots,m$, relates to how well the chosen SC-type estimators approximate the post-intervention counterfactual. This is the usual source of bias we encounter when using any SC-type estimator. The second, $B^{j}_{te},j=1,\ldots,m$, is induced by the fact that we are using units that are ”potentially affected” by the intervention in the donor pool. This type of bias solely depends on whether or not a given (”potentially affected”) unit receives weight and its respective treatment effect but not on how ”good” the chosen estimator is in recovering $Y_{jt}^{N},j=1,\ldots,m$.
@@
Assumption 2: As the number of pre-intervention periods $T_{0}$ goes to infinity, the SC-type used to estimate $\widehat{W}$ satisfy $\sum_{j=2}^{J}\widehat{w}_{j}{Y_{jt}^{N}}=Y_{1t}^{N}+o_{p}(1)$.
We state Assumption 2 in terms of estimated weights without any loss of generality and to simplify the discussion. Alternatively we can assume that the set of weights $\widehat{W}$ converges to a set of weights $W^{*}$ such that $\sum_{j=2}^{J}w^{*}_{j}{Y_{jt}^{N}}=Y_{1t}^{N}+o_{p}(1)$. When using the original SC-estimator, Assumption 2 amounts to assuming an (approximately) perfect fit. [^2] See the discussion in Abadie (2021), Ferman and Pinto (2021), and Powell (2018) about the lack of a perfect fit. This assumption can be relaxed by using an SC-type estimator that is robust to violation of the perfect fit assumption. We state assumption 2 in very general terms as our results apply to any SC-type estimator that can be written as a weighted average of the control units post-intervention outcomes. For example, for the original SCM one can use the results in Zhang et al. (2022), and state Assumption 2 accordingly. Depending on the chosen estimator, one can state Assumption 2 more precisely. A non-exhaustive list of possible estimators includes the one proposed in: Abadie and L’Hour 2021, Amjad et al. 2018, Ben-Michael et al. 2021, Ben-Michael et al. 2022, Doudchenko and Imbens 2017, Ferman and Pinto 2021, Kellogg et al. 2021, Xu 2017.
@@
It is important to notice that for each unit $j=2,\ldots,m$, if either $\gamma_{jt}$ or $\widehat{w}_{j}$ is zero, that unit does not induce any extra bias in $\widehat{\theta}_{1t}$. This implies that units that receive a low estimated weight need to have an extremely large effect to induce a non-negligible bias in $\widehat{\theta}_{1t}$. For this reason, units that receive a low weight can be relatively safely treated as ”pure controls” when estimating $\theta_{1t}$ in empirical applications. When using SC-type estimators that allow for negative weights one needs to check the magnitude of the weights given to ”potentially affected” units.
@@
For estimators that use a bias correction, one needs to subtract the estimated bias from the outcome of all units before running our iSCM (see equation (16) in Abadie 2021).
@@
Assumption 3: As the number of pre-intervention periods $T_{0}$ goes to infinity, the SC-estimator chosen to estimate $\widehat{L}^{i}$ satisfy $\widehat{Y}_{it}^{N}=Y_{it}^{N}+o_{p}(1),\ \forall\ i=2,\ldots,m.$
@@
This is a system of $m$ equations with $m$ unknowns, i.e., the treatment effect on the ”main treated” and the $m-1$ effects on the ”potentially affected” units.
@@
It is easy to show that $\widehat{\Omega}$ is always invertible, if $m<J-2$, except for some extreme cases. For example, $\widehat{\Omega}$ is not invertible if two units give weight 1 to each other and/or every single weight associated with the ”pure control” units is zero (see the appendix A.1 for more details). Note that, one can easily check the invertibility of $\widehat{\Omega}$ in the data.
@@
3 An example with a single ”potentially affected” unit
To further illustrate our results, it is useful to consider the special case in which, together with the ”main treated” unit, only one additional unit is ”potentially affected” by the intervention ($m=1$). First, we show the standard case and then the simplified case in which including the ”main treated” in the donor pool of the ”potentially affected” unit is not necessary.
With only one ”potentially affected” unit, the system of equation defined in Section 2, again ignoring the estimation errors, simplifies to